OpenStudy (anonymous):

Use the formula for the area of a rectangle A = lw to answer the question. A rectangle has a length of n units. Its width is 6 units less than its length. What is the area of the rectangle?

OpenStudy (bradely):

l = n b = n-6 A = lb =n(n-6) Source: http://www.mathskey.com/question2answer/